# Large Vision-Language Model Alignment and Misalignment: A Survey Through the Lens of Explainability

### v2

@@ 5 Conclusions
In this paper, we systematically survey alignment and misalignment in LVLMs through an explainability lens. We show that achieving proper alignment involves complex interactions between data quality, model architecture, and inference procedures. We categorize misalignment into object, attribute, and relational levels, providing a clear framework to understand these challenges and develop targeted solutions. Our review of mitigation strategies highlights approaches ranging from computationally intensive parameter-tuning methods to practical parameter-frozen solutions, each with trade-offs in effectiveness and feasibility. Lastly, we outline key future directions for building robust vision-language systems (see Appendix F), including development of standardized benchmarks, deeper investigation into explainability approaches to reveal internal alignment mechanisms, and architectural innovations to address fundamental limitations in current designs.
@@ Appendix A Overview of LVLM Architectures
Appendix B How to Measure Alignment?
@@
Appendix C More Reasons for Misalignment
@@
Appendix D Examples of LVLM Descriptions
POPE-Random POPE-Adversarial POPE-Popular Comp. Cost
Baseline
Mitigations
@@ Appendix D Examples of LVLM Descriptions
Appendix E Mitigation Methods Comparison
As shown in Table 3, we selected LLaVa-V1.5-7B as the baseline model and evaluated it on POPE benchmarks. We then tested three mitigation methods: SoM-LLaVA Yan et al. (2024), SID Huo et al. (2024), and LogicCheckGPT Wu et al. (2024a), to assess their effectiveness and efficiency in mitigating misalignment in the baseline model. The baseline model was loaded directly from Hugging Face, with parameters in float16 precision and all other implementation details set to default. Each mitigation method was implemented strictly according to its official GitHub repository, without modifications to core components, except for necessary adjustments such as path configurations.
We evaluated model performance using Accuracy, Precision, Recall, and F1-score, with the highest values underlined in the table. The results indicate that mitigation methods generally improve performance across most metrics compared to the baseline, particularly SoM-LLaVA and LogicCheckGPT, which achieved five and six highest scores, respectively. However, a deeper analysis reveals that some methods negatively impacted certain metrics. For instance, while LogicCheckGPT achieved the highest Precision on all benchmarks, it exhibited relatively lower Recall and F1 scores, sometimes even worse than the baseline. This occurs because LogicCheckGPT is conservative in predicting positives, meaning it only classifies a sample as positive when it is highly confident, resulting in fewer false positives but more false negatives. Interestingly, SoM-LLaVA consistently achieved the highest Recall across all benchmarks, indicating that it is less strict in determining positive cases compared to LogicCheckGPT. This highlights how different mitigation strategies lead to different prediction behaviors, emphasizing the need for future works to report performance across multiple metrics rather than focusing solely on a single high-performing metric.
@@ Appendix E Mitigation Methods Comparison
Appendix F Future Research Directions
@@ F.1 Standardized Benchmarks
F.2 Explainability based Diagnose

### v3

@@ 5 Conclusions
In this paper, we systematically survey alignment and misalignment in LVLMs through an explainability lens. We show that achieving proper alignment involves complex interactions between data quality, model architecture, and inference procedures. We categorize misalignment into object, attribute, and relational levels, providing a clear framework to understand these challenges and develop targeted solutions. Our review of mitigation strategies highlights approaches ranging from computationally intensive parameter-tuning methods to practical parameter-frozen solutions, each with trade-offs in effectiveness and feasibility. Lastly, we outline key future directions for building robust vision-language systems (see Appendix H).
@@
Acknowledgments
Mengnan Du is in part supported by National Science Foundation (NSF) Grant #2310261. Lu Cheng is supported by the National Science Foundation (NSF) Grant #2312862, NSF CAREER #2440542, NSF-Simons SkAI Institute, National Institutes of Health (NIH) #R01AG091762, Google Research Scholar Award, and a Cisco gift grant.
@@ Appendix A Overview of LVLM Architectures
A.1 How is Alignment Achieved?
The development of alignment in LVLMs progresses through three major stages (see Figure 3), each is built upon its predecessor to achieve increasingly sophisticated cross-modal integration. Additionally, we present prominent real-world LVLM architectures in Appendix A.
Stage 2: Adapter Fine-tuning.  The second stage involves fine-tuning an adapter module that bridges the pre-trained visual encoder with the language model Yang et al. (2024b). This stage introduces lightweight adapter architectures, which typically consist of simple components such as linear layers, MLPs, or cross-attention layers that learn to translate between visual and language model embedding spaces. For example, cross-attention layers can feed image encoder representations into the language model, enabling the model to attend to relevant visual features when generating text Meta AI (2024). A key characteristic of this approach is the preservation of the original capabilities of both the visual encoder and language model while learning to interface between them. During adapter training, while the visual encoder parameters may be updated, the language model parameters often remain frozen to maintain their original text capabilities. This intermediate stage is essential for establishing effective connections between modalities while preserving the specialized capabilities of each component.
Stage 3: End-to-End Fine-tuning.  The final stage involves comprehensive fine-tuning of the entire system, including the visual encoder, adapter, and LLM components together Zhai et al. (2024). This comprehensive approach allows for deeper integration and more sophisticated alignment between all components. It enables the model to learn task-specific optimizations that require coordinated adjustments across all modules. Through this process, the model develops more advanced cross-modal understanding capabilities and facilitates the emergence of emergent behaviors that arise from the deep integration of visual and textual processing. This stage results in the highest performance but requires careful balancing to avoid catastrophic forgetting or degradation of pre-existing capabilities.
Appendix B Representation Misalignment
At the representation level, alignment refers to how closely visual representations $v\in\mathcal{V}$ and textual representations $t\in\mathcal{T}$ are mapped within the shared embedding space $\mathcal{E}$ of an LVLM. Representation misalignment occurs when semantically corresponding visual and textual representations occupy distant regions in this space, resulting in low similarity. This misalignment can have significant downstream consequences. For example, if the embedding of an image containing an apple lies far from the embedding of the word "apple" and is instead closer to the embedding of "dog", the model may incorrectly interpret the image as depicting a dog. Such misalignment at the representation level undermines the model’s ability to ground visual input in the correct textual context. In this paper, we argue that representation misalignment fundamentally determines the upper ceiling of an LVLM’s performance, regardless of the quality of its language modeling or vision encoder in isolation.
Appendix C How to Measure Alignment?
@@
Appendix D Formal Theoretical Perspective
From a theoretical standpoint, visual and textual modalities can be modeled as distinct projections of a shared latent semantic space Ngiam et al. (2011). Let $Z\in\mathcal{Z}$ denote the latent variable representing the underlying world state or concept being observed. Images $x_{v}\in\mathcal{X}_{v}$ and text $x_{t}\in\mathcal{X}_{t}$ are generated via distinct, conditionally independent observation channels:
$x_{v}\sim p_{v}(x\mid Z),\quad x_{t}\sim p_{t}(x\mid Z).$ (1)
We assume visual and textual data are conditionally independent given the latent semantics:
$p(x_{v},x_{t}\mid Z)=p_{v}(x_{v}\mid Z)\cdot p_{t}(x_{t}\mid Z).$ (2)
This assumption reflects the intuitive idea that both images and texts encode complementary but overlapping information about the same underlying concept, such as an object, attribute, or relational.
Let $f_{v}:\mathcal{X}_{v}\rightarrow\mathbb{R}^{d}$ and $f_{t}:\mathcal{X}_{t}\rightarrow\mathbb{R}^{d}$ denote the learned encoders (e.g., a vision encoder and a LLM) that map raw inputs to $d$-dimensional embeddings. If both $f_{v}(x_{v})$ and $f_{t}(x_{t})$ preserve all the relevant information about the latent variable $Z$, they are said to be sufficient statistics for $Z$:
$I(Z;f_{v}(x_{v}))=I(Z;x_{v})$ (3)
$I(Z;f_{t}(x_{t}))=I(Z;x_{t}),$ (4)
where symbol $I(A;B)$ denotes mutual information between variables $A$ and $B$. This implies that no semantic information about $Z$ is lost in the representation process Tishby and Zaslavsky (2015); Poole et al. (2019). Consequently, the learned representations $f_{v}(x_{v})$ and $f_{t}(x_{t})$ should be functionally equivalent in terms of semantic content, as they reflect the same underlying meaning. This sets the foundation for representational alignment.
Appendix E More Reasons for Misalignment
@@
Appendix F Examples of LVLM Descriptions
@@
Appendix G Detailed Analysis of Mitigation Methods Comparison
As shown in Table 1, we evaluated model performance using Accuracy, Precision, Recall, and F1-score, with the highest values underlined in the table. The results indicate that mitigation methods generally improve performance across most metrics compared to the baseline, particularly SoM-LLaVA and LogicCheckGPT, which achieved five and six highest scores, respectively. However, a deeper analysis reveals that some methods negatively impacted certain metrics. For instance, while LogicCheckGPT achieved the highest Precision on all benchmarks, it exhibited relatively lower Recall and F1 scores, sometimes even worse than the baseline. This occurs because LogicCheckGPT is conservative in predicting positives, meaning it only classifies a sample as positive when it is highly confident, resulting in fewer false positives but more false negatives. Interestingly, SoM-LLaVA consistently achieved the highest Recall across all benchmarks, indicating that it is less strict in determining positive cases compared to LogicCheckGPT. This highlights how different mitigation strategies lead to different prediction behaviors, emphasizing the need for future works to report performance across multiple metrics rather than focusing solely on a single high-performing metric.
@@ Appendix G Detailed Analysis of Mitigation Methods Comparison
Appendix H Future Research Directions
@@ H.1 Standardized Benchmarks
H.2 Explainability based Diagnose
@@ H.3 Architectural Innovations
H.4 Real-world Implications
While LVLM misalignment mitigation strategies have achieved impressive improvements, most of these improvements are primarily validated on predefined benchmarks, with limited translation to real-world applications such as autonomous driving, medical imaging, or vision-focused assistive AI. However, different types of misalignment can have serious consequences in these domains. For instance, in autonomous driving, object or relational misalignment may cause the system to overlook pedestrians or misinterpret traffic scenarios, leading to safety-critical failures. In medical imaging, attribute misalignment could result in misdiagnosis by inaccurately describing clinical indicators. In assistive AI, especially for visually impaired users, relational or object misalignment may lead to incorrect environmental descriptions, potentially putting users at risk. These high-stakes scenarios underscore the urgent need to bridge the gap between benchmark-driven progress and deployment-focused validation. It is essential that future work moves beyond controlled datasets and begins testing and adapting misalignment mitigation methods in real-world settings to ensure their reliability, safety, and practical utility.